OpenStudy (anonymous):

Express as a single logarithm. Simplify, if possible. loga2x + 3(logax - logay) The a's are subscript

OpenStudy (anonymous):

\[\log_{a} 2x+3(\log_{a} x-\log_{a} y)=\log_{a} 2x+3\log_{a} \frac{ x }{y }=\log_{a} 2x+\log_{a} \left( \frac{ x }{ y } \right)^3 \] \[=\log_{a} 2x \frac{ x^3 }{ y^3 }=\log_{a} \frac{ 2x^4 }{y^3 }\]

OpenStudy (anonymous):

thank you so much surjithayer, wqe really appreciate your help